FERMI MIXING AND FORBIDDEN TRANSITIONS IN METHANOL-D_{1} AND CONFIRMATION OF OP-TICALLY PUMPED FAR-INFRARED LASER LINES

\underline{INDRANATH MUKHOPADHYAY}, Laser Programme, Centre for Advanced Technology, Indore 452 013, India.

This paper reports ”Forbidden” transitions in the Fourier transform infrared (FTIR) spectrum, involving highly excited torsional levels within the ground vibrational state for dK_{p}=0, K_{p}=4, of methanol- D_{1} (CH_{2}DOH) which become allowed through ”Fermi” interaction with the OCD-bending vibrational state. This is the first time such transitions have been detected for any asymmetric- asymmetric molecule. The present findings represent a breakthrough in the assignment processfor asymmetrically substituted methanol, providing ample data for modeling of the Hamiltonian of the molecule. The findings are consistent with the far-infrared (FIR) laser lines pumped by P(46) CO_{2} laser line of the 10.6 m band in CH_{2}DOH, which was earlier associated with an upshifted infrared (IR) transition but until now the mechanism of the anomalous shift remained a mystery. The results will be discussed in terms of the ”Fermi” interaction and closed combination loops. Comments will be made on the internal rotational potential energy coefficients.